I need help with this problem

Use the variation of parameters X(t) = A x CD + FC A per the solution and anrwer are not shown
Answer #1

© A[=>] i freno ( el i flere lot Y = AX+ fits by Variation of Parameter X=18)* + Clemens det (A-13) = 1 8-1 -4.= –6470 14 -for 1,6 How they :-:11%)=(:1 PPPELPPPPPPPP! - if we. Take Vest . then Vicat * = 65)+ C[i]ent let = $(t) ūlti with Bitl= ( 18nowa pole, Pies - et me (*) *** 71 *** t *[**] - [6.54) ( 5 Rest z filetat 1 = ( 6 Eilgts-el + C) es constant On Integrate. oNow x= c/187+ 6 (3) est Xects ☺ where Xylt as given in